OpenStudy (anonymous):

simplify: e^3lnx..plzz explain how u got ur answer

OpenStudy (anonymous):

not clear what this is saying. it is \[e^{3\ln(x)}\]?

OpenStudy (anonymous):

if that is what you meant, then it is \[e^{3\ln(x)}=e^{\ln(x^3)}=x^3\] the first

OpenStudy (anonymous):

the first equality because \[n\ln(x)=\ln(x^n)\] and the second because log and exponential are inverse functions

OpenStudy (anonymous):

teh qustion is: \[e ^{3lnx} \]

OpenStudy (anonymous):

then the answer i wrote is correct

OpenStudy (anonymous):

i just wrote \[\ln(x)\] instead of the more lazy \[\ln x\]